# Inspection-Guided Randomization: A Flexible and Transparent Restricted Randomization Framework for Better Experimental Design

### v1

@@ 3.4 Analyzing IGR experiments: Randomization inference
We take a design-based approach to inference, meaning we view the potential outcomes as fixed conditional on the study sample, so that the only randomness in the study arises through the randomness in the treatment assignment mechanism. In IGR specifically, randomness arises in Step 6. Randomization, where the observed allocation is sampled from the restricted assignment mechanism, $P_{IGR}$. Randomization then forms the basis for inference, and we can perform a Fisher randomization test to evaluate null hypotheses. Specifically, we compute the value of the test statistic under the null hypothesis for different possible assignment vectors. With IGR, the empirical null distribution for the test statistic is derived over the set of assignment vectors $\mathbf{z}^{j}$ in $\mathcal{Z}_{pool}$ that have $p_{\mathbf{z}_{j}}>0$. We compare the observed value of the test statistic for $\mathbf{z}^{obs}$ to the empirically derived null distribution of the test statistic to obtain a p-value.
4 Applying IGR: Two Vignettes
In this section, we use vignettes and simulated experiments to illustrate how IGR can be used to achieve design desiderata in two distinct experimental settings:
  1. 1.
    Group formation experiments. In a group formation experiment, the challenge is to minimize differences in most baseline covariates, while creating large contrasts in a salient covariate of interest. Individuals are randomized into different groups, typically to construct groups of certain compositions that are believed to have different responses to the intervention (Basse et al.,, 2024). For example, a group may be the set of college freshmen assigned to the same dorm room, and composition may be the number of individuals in a room who were “high academic achievers” in high school (Sacerdote,, 2001). We consider group formation trials where composition is defined in terms of a categorical salient attribute (e.g., high or low prior academic achievement) and is count-based (e.g., number of individuals with high academic achievement). A well-designed experiment in this setting has groups balanced on all observed covariates other than the salient attribute itself.
  2. 2.
    Experiments with interference. Interference occurs when the treatment assigned to one observational unit affects the outcome of a different unit. With many interventions, interference is expected. For example, consider a behavioral intervention that teaches participants new information or skills. Participants in the treatment arm who receive the treatment may share what they have learned with their friends, some of whom may be assigned to the control arm. A common target estimand in experiments with interference is the total average treatment effect, or the difference in outcomes if everyone were to receive the treatment compared to if everyone were to receive the control. A well-designed experiment in this setting assigns treatments such that those in the treatment (control) arm experience a universe that is as close as possible to an all-treated (all-control) universe while also balancing covariates across arms.
In each vignette, we simulate an experiment designed using IGR and benchmark the results against alternative randomization designs (see Appendix A for the general simulation pipeline). Code is available at https://github.com/wangmagg/inspection-guided-randomization.
4.1 Vignette 1: IGR for group formation experiments
In this first vignette, we illustrate how to use IGR in the design of a group formation experiment where we are interested in evaluating how group composition influences the effect of a group-level intervention. Various kinds of inter-group comparisons may be of interest, including between treated and control groups with the same compositions (to target the treatment effect under a fixed composition) and between treated and treated groups with different compositions (to target the effect of the composition under a fixed treatment). In addition, we may also want to make difference-in-differences comparisons between the treatment effect under one composition and the treatment effect under a different composition.
The design goals in this setting are two-fold. First, we would like to sort individuals into groups such that certain pre-specified group compositions are realized. We focus in particular on the case where composition is characterized by the proportion of the group that possesses a particular salient attribute. Second, we would like to maintain balance on all covariates other than those used to define composition. A key point of this vignette is to highlight that in the Enumeration step of IGR, we can use different methods to sample allocations. Here, we sample from a blocking-based group formation randomization assignment mechanism.
We first discuss the choices we make in the Specification and Enumeration steps to meet these goals, then present results from a simulated group formation experiment.
4.1.1 Specification and Enumeration for a group formation experiment
We begin by establishing notation and formally clarifying the design goals in a group formation experiment. Suppose we want to form $N_{G}$ total groups. Let $g_{i}\in\{1,\dots N_{G}\}$ denote the group assignment for individual $i$ and $\mathbf{g}$ denote the $N$-vector of group assignments. Let $a$ be the salient attribute used to define group composition. Let $x_{ia}=1$ for individuals who have this attribute and $x_{ia}=0$ for individuals who do not have this attribute. We then define the composition of group $h$, denoted $\rho_{a}(h)$, as the proportion of the group that possess attribute $a$, where
$\rho_{a}(\mathbf{g},h,\mathbf{X})=\frac{\sum_{i=1}^{N}\mathbbm{1}\{x_{ia}=1\}\mathbbm{1}\{g_{i}=h\}}{\sum_{i=1}^{N}\mathbbm{1}\{g_{i}=h\}}$ (4.1.1)
Suppose that we would like to form groups that have $L$ unique compositions, $\mathcal{C}=\{c_{a}^{1},c_{a}^{2},\dots c_{a}^{L}\}$. Furthermore, we seek to form at least two groups with each composition, such that, for every composition, there can be at least one group with that composition assigned to treatment and one assigned to control.
Specification
We specify the target effect estimands as the average treatment effects under particular group compositions of interest. We denote the potential outcome for individual $i$ in group $h$ under treatment $z_{h}$ and composition $c_{a}^{l}\in\mathcal{C}$ as $Y_{ih}(z_{h},c_{a}^{l})$. The target estimands are $\tau_{l}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(1,c_{a}^{l})-Y_{i}(0,c_{a}^{l})\ \text{for all}\ c_{a}^{l}\in\mathcal{C}$. We specify one inspection metric that inspects allocations for balance and one for whether desired group compositions are realized.
Inspection Metric 1 (Balance): Standardized mean difference and Mahalanobis distance are commonly used balance metrics in two-arm studies. Here, we extend these metrics to the group formation experiment setting. Let $\mathbf{z}\in\{0,1\}^{N_{G}}$ denote the vector of treatment assignments across groups. Since there are both random treatment assignments $\mathbf{z}$ and random group assignments $\mathbf{g}$, we use fitness functions of the form $f(\mathbf{z},\mathbf{g},\mathbf{X})$.
We define $m^{(\texttt{SumMaxAbsSMD-G})}$ and $m^{(\texttt{MaxMahalanobis-G})}$ as
$m^{(\texttt{SumMaxAbsSMD-G})}(\mathbf{z},\mathbf{g},\mathbf{X})$ $=\sum_{j=1,j\neq a}^{p}\max_{(h_{1},h_{2})\in\mathcal{G}\times\mathcal{G}}\frac{\left|\bar{X}_{\cdot j}^{(h_{1})}-\bar{X}_{\cdot j}^{(h_{2})}\right|}{\sqrt{\left(s_{\cdot j}^{(h_{1})}\right)^{2}+\left(s_{\cdot j}^{(h_{2})}\right)^{2}}}$ (4.1.2)
$m^{(\texttt{MaxMahalanobis-G})}(\mathbf{z},\mathbf{g},\mathbf{X})$ $=\max_{(h_{1},h_{2})\in\mathcal{G}\times\mathcal{G}}\left(\bar{\mathbf{X}}_{\setminus a}^{(h_{1})}-\bar{\mathbf{X}}_{\setminus a}^{(h_{2})}\right)S_{\setminus a}^{-1}\left(\bar{\mathbf{X}}_{\setminus a}^{(h_{1})}-\bar{\mathbf{X}}_{\setminus a}^{(h_{2})}\right)^{T}$ (4.1.3)
where $\mathcal{G}=\{0,1,2,\cdots,N_{G}-1\}$ is the set of group indices, $\bar{X}_{\cdot j}^{(g)}$ and $s_{\cdot j}^{(h)}$ are the mean and standard deviation of covariate $j$ in group $h$, $\mathbf{\bar{X}}_{\setminus a}^{(h)}$ is the vector of covariate means in group $h$ excluding the salient attribute $a$, and $S_{\setminus a}$ is the sample covariance matrix excluding $a$.
Inspection Metric 2 (Composition Realization): To ensure allocations have the desired group compositions, we define the metric $m^{(\texttt{DesiredComps})}$ as
$m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=\begin{cases}1&\rho_{a}(\mathbf{g},h,\mathbf{X})\in\mathcal{C}\ \text{for all}\ h\in\mathcal{G},\\ &\mathbbm{1}\{\gamma^{0}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})>0\}\cdot\mathbbm{1}\{\gamma^{1}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})>0\}=1\\ &\ \ \ \ \text{for all}\ c_{a}^{l}\in\mathcal{C}\\ 0&\text{otherwise}\end{cases}$ (4.1.4)
where $\gamma^{0}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})=\sum_{h\in\mathcal{G}}\mathbbm{1}\{\rho_{a}(\mathbf{g},h,\mathbf{X})=c_{a}^{l}\}\cdot(1-\mathbf{z_{h}})$ is the number of control groups with composition $c_{a}^{l}$ and $\gamma^{1}(\mathbf{z},\mathbf{g},\mathbf{X},c_{a}^{l})=\sum_{h\in\mathcal{G}}\mathbbm{1}\{\rho_{a}(\mathbf{g},h,\mathbf{X})=c_{a}^{l}\}\cdot\mathbf{z_{h}}$ is the number of treated groups with composition $c_{a}^{l}$.
Combining these metrics, we define the fitness function
$\begin{split}f^{(\texttt{SumMaxAbsSMD-GD})}(\mathbf{z},\mathbf{g},\mathbf{X})&=\Theta(m^{(\texttt{SumMaxAbsSMD-G})}(\mathbf{z},\mathbf{g},\mathbf{X}),m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X}))\\ &=\begin{cases}m^{(\texttt{SumMaxAbsSMD-G})}(\mathbf{z},\mathbf{g},\mathbf{X}))&m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=1\\ \infty&m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})=0\end{cases}\end{split}$ (4.1.5)
The fitness function $f^{(\texttt{MaxMahalanobis-GD})}$ is defined analogously. Lastly, we specify the restriction rule $r(s)=0,s\geq s^{*}$, $r(s)=1,s<s^{*}$, where $s^{*}$ is the $(100\cdot\frac{m}{M})^{\text{th}}$-percentile of scores. By applying this rule, $m$ out of the $M$ enumerated allocations are accepted.
Enumeration
For simplicity, we assume that we want to form groups of equal sizes such that there are exactly two groups with each composition, one of which is assigned treatment and one assigned control. We furthermore assume that the number of individuals that possess the salient attribute $a$ permits forming the desired groups. Then, to enumerate a candidate allocation, we first split the study population into two blocks, $\mathcal{B}_{1}=\{i:X_{ia}=1\}$ and $\mathcal{B}_{2}=\{i:X_{ia}=0\}$ We sample $c_{a}^{j}\cdot\frac{N}{N_{G}}$ individuals from $\mathcal{B}_{1}$ and $(1-c_{a}^{j})\cdot\frac{N}{N_{G}}$ from $\mathcal{B}_{2}$ uniformly at random, without replacement, to place into each pair of groups with the composition $c_{a}^{j}$. For every pair of groups with the same composition, we assign one group to treatment and one group to control. Under this group formation randomization assignment mechanism, $m^{(\texttt{DesiredComps})}(\mathbf{z},\mathbf{g},\mathbf{X})$ is guaranteed to equal 1. We refer to this assignment mechanism as group formation randomization (GFR). Note that, if we were to instead sample candidate allocations from an assignment mechanism that is agnostic to group composition, it would require enumerating an extremely large number of allocations to obtain any allocations that have the desired group compositions.
4.1.2 Simulated group formation experiment with perceived gender stereotyping as a contextual moderator
Data generation
Our group formation experiment simulation is inspired by stereotype threat interventions that have shown heterogeneous effects based on the gender composition of a group, where composition determines whether or not the individuals in the group perceive the presence of gender stereotyping (Grover et al.,, 2017; Carroll et al.,, 2023). The data generation process is modeled based on an unpublished educational intervention study that tested how different types of feedback on class assignments affect students’ grades. In the first two weeks of the class, the researchers collected information about students (age, gender, and major) and early course assignments (scores on a homework). After the second week of class, the researchers aimed to assign all $N$ students in the class to $k$ treatment arms (feedback conditions) with similar observed characteristics.
We simulate the study data set to encode theory-based covariance between student characteristics (age, gender, and college major), latent traits (ability and confidence), and course performance (homework score and final exam grade). The encoded covariance structure creates conditions under which it is difficult to achieve balanced assignment to multiple groups. We let the salient attribute $a$ be the gender covariate, where $x_{ia}=1$ if individual $i$ is a man, and we define group composition $\rho_{a}(g)$ as the proportion of individuals in group $g$ that are men. We designate exam grade as the outcome variable. To generate the observed outcomes, we apply an additive treatment effect, with the magnitude of the treatment effect varying across compositions.
We perform simulations with $M=100,000$ enumerated allocations and $m=500$ accepted allocations, which is in line with the number of enumerated and accepted allocations in prior work (Li et al.,, 2016; Watson et al.,, 2021; Nordin and Schultzberg,, 2022). We compare applying IGR using both a balance and a composition realization metric to applying IGR using a composition realization metric only. The latter design is functionally equivalent to sampling from the group formation randomization assignment mechanism described above, so we simply refer to it as “GFR” in the results. Further details on the data-generating process and the parameter settings used in the simulations can be found in Appendix B.
Results
Using the fitness function $f^{(\texttt{SumMaxAbsSMD-GD})}$ or $f^{(\texttt{MaxMahalanobis-GD})}$ (4.1.2) in conjunction with a thresholding restriction rule passes visual inspections for the “Discriminatory Power” and “Over-restriction” checks in the Evaluation & Adaptation step of IGR (Figure 4.1.1). By using $f^{(\texttt{SumMaxAbsSMD-GD})}$ or $f^{(\texttt{MaxMahalanobis-GD})}$ (4.1.5) as the fitness function, only allocations that have the desired compositions are accepted, and those with large imbalance on covariates are filtered out (Figure 4.1.2). In contrast, when individuals are randomized into groups of particular compositions using group formation randomization only, i.e. we apply the inspection metric $m^{(\texttt{DesiredComps})}$ as the fitness function, covariates are less balanced. Notably, including a balance metric helps improve balance on “Homework”, which is correlated with the salient attribute,“Gender”, and thus is highly imbalanced under GFR due to the differing gender compositions across groups. In Table 4.1.1, we show that, as expected, including a balance metric in IGR reduces the root mean-squared error (RMSE) of treatment effect estimates compared to using a composition metric alone.
Figure 4.1.1: [Vignette 1, $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\},\ N=120$], Visualizations for the Evaluation & Adaptation step: discriminatory power and over-restriction. (top) For both $m^{(\texttt{SumMaxAbsSMD-GD})}$ and $m^{(\texttt{MaxMahalanobis-GD})}$, there is substantive differentiation in score across candidate allocations, both when the allocations are enumerated without ($\mathcal{Z}_{pool}$) and with the addition of genetic algorithms ($\mathcal{Z}^{*}_{pool}$). $\mathcal{Z}^{*}_{pool}$ excludes the worst-scoring allocations in the right tail of $\mathcal{Z}_{pool}$. (bottom) IGR and IGRg only slightly increase pairwise assignment correlation relative to GFR.
Figure 4.1.2: [Vignette 1, $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\},\ N=120$] Standardized mean difference across all pairwise group comparisons for accepted allocations under Group Formation Randomization (GFR) and IGR/IGRg. Groups 0 and 1 are 50 % men, 2 and 3 are 30% men, and 4 and 5 are 70 % men. “Max” refers to the maximum SMD across all pairwise group comparisons. IGR improves balance on the non-salient covariates; using genetic algorithms (IGRg) does not noticeably enhance these improvements. Since “Homework” is correlated with “Gender”, it exhibits imbalance under GFR; this imbalance is reduced by including a balance metric in IGR. “Ability” and “Confidence” are latent covariates and are not included as direct inputs to the IGR balance metric. IGR improves balance on “Ability”, even though it is unobserved, because it is tightly correlated with “Homework.” To allow for a wider genetic algorithm search in IGRg, we permitted group sizes to range from 19 to 21 individuals and relaxed the composition realization constraint for allocations in child generations.
Group Comparison 0 vs 1 $(c_{a}^{1}=0.5,\text{ effect size }d=0.3)$ 2 vs 3 $(c_{a}^{2}=0.3,\text{ effect size }d=0.5)$ 4 vs 5 $(c_{a}^{3}=0.7,\text{ effect size d }=0.1)$ RMSE %$\%\text{RMSE}_{GFR}$ Rej. Rate RMSE %$\%\text{RMSE}_{GFR}$ Rej. Rate RMSE %$\%\text{RMSE}_{GFR}$ Rej. Rate Design Fitness GFR 0.39$\pm$0.02 Ref 0.16$\pm$0.02 0.40$\pm$0.02 Ref 0.35$\pm$0.04 0.39$\pm$0.02 Ref 0.05$\pm$0.00 IGR MaxMahalanobis-GD 58.49$\pm$2.57 0.37$\pm$0.04 57.07$\pm$3.68 0.78$\pm$0.02 55.91$\pm$4.44 0.08$\pm$0.01 SumMaxAbsSMD-GD 49.90$\pm$3.60 0.45$\pm$0.02 48.41$\pm$3.32 0.89$\pm$0.03 48.21$\pm$1.50 0.08$\pm$0.01 IGRg MaxMahalanobis-GD 56.82$\pm$2.97 0.38$\pm$0.04 55.52$\pm$4.41 0.80$\pm$0.04 56.39$\pm$3.69 0.08$\pm$0.01 SumMaxAbsSMD-GD 48.83$\pm$2.69 0.46$\pm$0.03 47.61$\pm$3.91 0.90$\pm$0.04 47.82$\pm$3.03 0.08$\pm$0.01
Table 4.1.1: [Vignette 1, $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\},\ N=120$] Root mean-squared error (RMSE) in treatment effect estimates and rejection rate of sharp null for IGR/IGRg compared to GFR in a group formation experiment. IGR/IGRg with both a balance and a composition realization metric reduces RMSE across group comparisons compared to GFR (equivalently, IGR with a composition metric alone). Means and standard deviations are computed over 5 data samples.
4.2 Vignette 2: IGR for experiments with interference
In the second vignette, we illustrate how to use IGR for an experiment where we expect interference, which occurs when one individual’s treatment affects a different individual’s outcome (Rosenbaum,, 2007). Interference is often present with behavioral interventions, where information can diffuse or spill over along social ties (Valente,, 2012; Cai et al.,, 2015; Kim et al.,, 2015). When interference is present, a common target treatment effect estimand is the total average treatment effect, or a comparison of the average outcome if everyone were to receive treatment versus if everyone were to receive control (Hudgens and Halloran,, 2008). Failing to properly account for interference can make estimates of the total treatment effect both biased and imprecise (Eckles et al.,, 2017). This in turn can result in misleading conclusions (e.g., false nulls) and misinformed recommendations for policy-making (e.g., failure to implement an effective intervention).
A common approach to controlling interference is to use cluster randomization, where randomization is done on the cluster level such that all individuals in the same cluster receive the same treatment (Hayes and Moulton,, 2017). Given enough separation between clusters, there is likely no spillover between different treatment arms, and comparing treated to control clusters can give an unbiased estimate of the total average treatment effect. However, it may sometimes be difficult to identify clusters that are sufficiently separated to avoid any inter-cluster interference. In other cases, it may be possible to identify a few large clusters that are well-separated, such as villages that are far apart geographically. However, randomizing only a few clusters leads to low power, especially if there is strong intra-cluster correlation on prognostic covariates. Often, for clusters that form organically, individuals within a cluster are more similar than those in different clusters, so intra-cluster correlation is expected.
Within the IGR design framework, we can be more flexible with the way that we control interference. Leveraging prior knowledge, which may not necessarily be complete or fully correct, we can use an interference-targeting fitness function to control the amount of anticipated interference in a more precise way than with cluster randomization. Even if we do not have perfect knowledge of the interference structure (e.g., which individuals interact to produce spillover), we can still develop “coarse” metrics that measure interference.
In this section, we discuss the choices we make in the Specification and Enumeration steps of IGR, then present simulations based on a sexual assault prevention intervention.
4.2.1 Specification and Enumeration for an experiment with interference
Specification
We define the target effect estimand to be the total average treatment effect, $\tau_{TATE}=\frac{1}{N}\sum_{i=1}^{N}Y_{i}(\mathbf{1})-Y_{i}(\mathbf{0})$ where $\mathbf{1}$ is the all-ones vector (everyone is treated) and $\mathbf{0}$ is the all-zeros vector (everyone is control). We specify one inspection metric for interference and one for balance.
Inspection Metric 1 (Interference): We present two example metrics for controlling interference. The first can be used even with coarse knowledge of the interference structure whereas the second requires access to the interference network.
$m^{(\texttt{InvMinEuclideanDist})}=\left(\min_{(i,j)\in[N]\times[N],z_{i}\neq z_{j}}||\mathbf{l}_{i}-\mathbf{l}_{j}||_{2}^{2}\right)^{-1}$ (4.2.1)
where $z_{i},z_{j}\in\{0,1\}$ are the treatment assignments for units $i$ and $j$, respectively and $\mathbf{l}_{i},\mathbf{l}_{j}\in\mathbb{R}^{2}$ are coordinates for $i$ and $j$. While we use Euclidean distance here, other distance functions can also be used. It is also possible to expand upon a simple distance function by incorporating known information about ease of travel, such as the presence or absence of interconnecting roads.
In some cases, we may have more detailed knowledge of a network of interactions between individuals. For instance, we can ask participants to list the friends with whom they communicate regularly and build a network based on each participant’s list of contacts (Cai et al.,, 2015; Kim et al.,, 2015). In the network interference literature, it is common to assume that an individual is “exposed” through the treatments of their neighbors in the network via a neighborhood exposure model (Aronow and Samii,, 2017; Manski,, 2013).
Suppose we have a priori knowledge of the underlying network and seek to limit the number of control individuals exposed to treatment. This motivates an inspection metric that measures the number of control individuals that are exposed to treatment under a particular allocation. Let $\mathbf{A}$ be an $N\times N$ adjacency matrix such that $A_{ij}=1$ if $i$ and $j$ are joined by an edge in the network and $A_{ij}=0$ otherwise. Let $\mathcal{N}(i)$ denote the neighborhood of $i$ and let $\mathbf{z}_{\mathcal{N}(i)}$ denote the treatment assignments for the individuals in this neighborhood. Let $h(\cdot)$ be an exposure function where $h(\mathbf{z}_{\mathcal{N}(i)})=1$ if individual $i$ is exposed to the treatment and $h(\mathbf{z}_{\mathcal{N}(i)})=0$ otherwise. For example, for a “one-neighbor” exposure model, $h(\mathbf{z}_{\mathcal{N}(i)})=\mathbbm{1}\{\sum_{j\in\mathcal{N}(i)}z_{j}>0\}$. Using this notation, we define the inspection metric $m^{(\texttt{FracCtrlExposed})}$ as
$m^{(\texttt{FracCtrlExposed})}(\mathbf{z},\mathbf{A})=\frac{\sum_{i=1}^{N}\mathbbm{1}\{h(\mathbf{z}_{\mathcal{N}(i)})\}\cdot(1-z_{i})}{\sum_{i=1}^{N}(1-z_{i})}$ (4.2.2)
Inspection Metric 2 (Balance): We assess covariate balance using the metric $m^{(\texttt{MaxMahalanobis})}$, which is defined similarly to $m^{(\texttt{MaxMahalanobis-G})}$ (4.1.2) except there is no salient attribute that is excluded.
Combining the interference and balance metrics, we define the fitness function
$\begin{split}f^{(\texttt{F+M})}(\mathbf{z},\mathbf{X})&=\Theta(m^{(\texttt{FracControlExposed})}(\mathbf{z},\mathbf{X}),m^{(\texttt{MaxMahalanobis})}(\mathbf{z},\mathbf{X}))\\ &=w_{1}\cdot m^{(\texttt{FracControlExposed})}(\mathbf{z},\mathbf{X})+w_{2}\cdot m^{(\texttt{MaxMahalanobis})}\cdot(\mathbf{z},\mathbf{X})\end{split}$ (4.2.3)
where $w_{1}$ and $w_{2}$ are scalar weights that can be used to tune whether there is heavier emphasis on the balance desideratum or the interference desideratum. The fitness function $f^{(\texttt{I+M})}$ is defined analogously, replacing $m^{(\texttt{FracControlExposed})}$ with $m^{(\texttt{InvMinEuclideanDist})}$. Lastly, we use the thresholding restriction rule as defined in Vignette 1 (4.1).
Enumeration
We sample from a complete randomization assignment mechanism that assigns an equal number of units to the treatment and control arms.
4.2.2 Simulated experiment with interference inspired by a sexual assault prevention intervention
Data generation
Results
In this vignette, as was the case in Vignette 1 (Section 4.1), there are design desiderata that are in tension with one another — minimizing interference and maximizing balance. With IGR, the study analyst can leverage visualizations similar to those we offer in Figure 4.2.1 as part of the Evaluation & Adaptation step to gain insights into the nature of these types of desiderata trade-offs, which will likely vary by study. Even in our simple simulation, varying the parameter settings for the data-generating process produces unique trade-offs between balance and interference desiderata. Specifically, when the probability that two individuals are connected decays more strongly with the distance between their respective schools, i.e. the interference network is sparser, the trade-off between balance and interference becomes less pronounced (Figure 4.2.1).
We show through visualizations that the “Discriminatory Power” and “Over-restriction” checks in the Evaluation & Adaptation step are satisfied when we use $f^{(\texttt{F+M})}$ (4.2.3) as fitness function in conjunction with a thresholding restriction rule (Figure 4.2.2, top and bottom rows). We furthermore show that modifying how the interference and balance metrics are aggregated can be an effective means of tuning desiderata prioritization. By adjusting the weights $w_{1}$ and $w_{2}$ assigned to $m^{(\texttt{FracControlExposed})}$ and $m^{(\texttt{MaxMahalanobis})}$, respectively, we can prioritize accepting allocations with better interference properties or with better balance properties (Figure 4.2.2, middle row); achieving both simultaneously is difficult due to the trade-offs depicted in Figure 4.2.1. Without access to the underlying interference network, we can use a Euclidean distance-based proxy metric, $m^{(\texttt{InvMinEuclideanDist})}$ to measure interference. While $f^{(\texttt{I+M})}$ filters out a subset of high-interference allocations, it is not as effective as $f^{(\texttt{F+M})}$ (Appendix Figure D.1). Finally, prioritizing interference reduction in IGR decreases bias, particularly in denser graphs, but also generally leads to variance inflation (Table 4.2.1).
Figure 4.2.1: [Vignette 2, $N=4000$] Visualization of balance and interference desiderata trade-offs in candidate allocations. Each bin in the heatmap represents a count of the number of candidate allocations with $m^{(\texttt{MaxMahalanobis})}$ and $m^{(\texttt{FracControlExposed})}$ values that fall into that bin. Darker colors indicate a higher density of allocations. As $\gamma$ increases, the probability two individuals are connected decays more strongly with increasing Euclidean distance between the individuals’ respective schools. A downward-sloping trend is suggestive of a trade-off between achieving balance (small values on the x-axis) and reducing interference (small values on the y-axis). We observe a stronger trade-off for denser graphs.
Figure 4.2.2: [Vignette 2, sparsity $\gamma=0.5$, $N=4000$] Visualizations for the Evaluation & Adaptation step when $f^{(\texttt{F+M})}$ is used as fitness function. (top) For all three weighting schemes, there is differentiation in $f(\mathbf{z})$ across candidate allocations, whether they are enumerated naively ($\mathcal{Z}_{pool}$) or with the use of genetic algorithms ($\mathcal{Z}^{*}_{pool}$). (middle) Each point represents a single accepted allocation. Complete randomization (CR) is plotted for reference. When prioritizing interference reduction, we accept more imbalanced allocations. When prioritizing balance, we are less effective in reducing interference. The FracExpo score distribution is bimodal because FracExpo is asymmetric and tends to be larger for mirror allocations. (bottom) IGR/IGRg increases pairwise assignment correlation slightly when interference is prioritized.
Design Fitness Bias %$\%\text{Bias}_{CR}$ Var %$\%\text{Var}_{CR}$ Rej. Rate CR -0.41$\pm$0.14 Ref 0.61$\pm$0.10 Ref 0.15$\pm$0.04 IGR 0.25*MaxMahalanobis + 0.75*FracExpo 87.88$\pm$0.77 140.61$\pm$26.80 0.13$\pm$0.04 0.50*MaxMahalanobis + 0.50*FracExpo 93.99$\pm$0.42 22.35$\pm$3.15 0.55$\pm$0.17 0.75*MaxMahalanobis + 0.25*FracExpo 99.14$\pm$0.67 4.49$\pm$0.20 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 99.65$\pm$0.83 33.24$\pm$21.05 0.49$\pm$0.37 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 100.95$\pm$0.27 15.00$\pm$8.20 0.72$\pm$0.27 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 101.66$\pm$0.29 7.83$\pm$2.30 0.91$\pm$0.08 IGRg 0.25*MaxMahalanobis + 0.75*FracExpo 88.19$\pm$1.07 113.05$\pm$26.97 0.13$\pm$0.06 0.50*MaxMahalanobis + 0.50*FracExpo 96.21$\pm$0.87 9.36$\pm$1.53 0.88$\pm$0.10 0.75*MaxMahalanobis + 0.25*FracExpo 99.80$\pm$0.82 2.62$\pm$0.21 1.00$\pm$0.00 0.25*MaxMahalanobis + 0.75*InvMinEuclidDist 99.50$\pm$1.09 30.26$\pm$21.27 0.56$\pm$0.37 0.50*MaxMahalanobis + 0.50*InvMinEuclidDist 100.90$\pm$0.32 7.05$\pm$4.61 0.89$\pm$0.14 0.75*MaxMahalanobis + 0.25*InvMinEuclidDist 101.83$\pm$0.34 2.58$\pm$0.71 1.00$\pm$0.00
[…]

### v2

@@ 5.2 Randomization Inference for IGR-Designed Experiments
We take a design-based approach to inference, meaning we view the potential outcomes as fixed conditional on the study sample, so that the only randomness in the study arises through the randomness in the treatment assignment mechanism. In IGR specifically, randomness comes from the Randomization step, where the observed allocation is sampled from the restricted assignment mechanism, $P_{IGR}$. Randomization then forms the basis for inference, and we can perform a Fisher randomization test to evaluate null hypotheses. Specifically, we compute the value of the test statistic under the null hypothesis for different possible assignment vectors. With IGR, the empirical null distribution for the test statistic is derived over the set of accepted assignment vectors $\mathbf{z}^{j}$ in $\mathcal{Z}_{pool}$ that have $p_{\mathbf{z}_{j}}>0$. We compare the observed value of the test statistic for $\mathbf{z}^{obs}$ to the empirically derived null distribution of the test statistic to obtain a p-value.
6 How IGR Improves Effect Estimates
Using IGR to design an experiment can lead to less biased, more precise effect estimates. We demonstrate these benefits with results from our simulated experiment (refer to Section 4.1 for a description of the simulation). In the Evaluation & Adaptation step in Section 4.2, we chose the fitness function, $f^{(\texttt{F+M})}(\mathbf{z},\mathbf{X})=0.75\cdot m^{(\texttt{FracExpo})}(\mathbf{z},\mathbf{X})+0.25\cdot m^{(\texttt{Mahalanobis})}(\mathbf{z},\mathbf{X})$, to apply to the design of the simulated experiment. Because the interference metric $m^{(\texttt{FracExpo})}$ is asymmetric, the resulting restricted assignment mechanism does not generally satisfy the mirror property unless we incorporate all mirror allocations, including those that would not be accepted by the restriction rule (refer to Section 5.1 for a discussion on the mirror property and estimation bias). To understand the implications of satisfying versus violating the mirror property, we examine effect estimates both with and without the inclusion of mirror allocations. We additionally assess how the estimation properties change when we apply different weights to the interference and balance metrics in the fitness function; below, we refer to the weight applied to the interference metric, $m^{(\texttt{FracExpo})}$, as $w_{Interference}$. We also show how, when the interference metric is upweighted in the design phase, the incurred covariate imbalance can be corrected for in the analysis phase through covariate adjustment. Results are reported based on five repeatedly drawn samples of the covariate and outcome data.
When mirrors are included and $w_{Interference}=0.75$, IGR reduces the absolute bias of the difference-in-means estimate ($88.31\pm 0.56\%$ of the absolute bias of an unrestricted design)(Figure 4). As the magnitude of the weight applied to the interference metric increases, we observe more bias reduction ($86.18\pm 0.37\%$ of the absolute bias of an unrestricted design when $w_{Interference}=1$). At small values of $w_{Interference}$, IGR decreases the variance of the difference-in-means estimate (e.g., variance at $w_{Interference}=0.125$ is $0.003\pm 0.0007$ compared to $0.61\pm 0.10$ for an unrestricted design). Increasing $w_{Interference}$ leads to variance inflation (e.g. variance at $w_{Interference}=0.875$ is $1.01\pm 0.28$). When $w_{Interference}=0.75$, leaving mirrors out leads to a large reduction in absolute bias for some data samples (e.g., $26.67\%$ of the absolute bias of the unrestricted design) and no reduction for others (e.g., $103.89\%$ of the absolute bias of the unrestricted design). This suggests that, in our example, including mirror allocations helps to make the estimation bias less dependent on the particular sample of individuals in the study.
Instead of a difference-in-means estimator, we can also fit a linear regression model that adjusts for observed baseline covariates. As mentioned in the Evaluation & Adaptation step in Section 4.2, using this estimator can be helpful when we choose to upweight the interference metric and downweight the balance metric in the design phase. As expected, the regression estimator substantially reduces variance compared to the difference-in-means estimator, particularly when $w_{Interference}$ is large (Figure 5). When mirror allocations are excluded, the bias in the regression estimate is more stable across data samples compared to the difference-in-means estimate. On the other hand, when mirror allocations are included, the regression estimate generally provides less bias reduction than the difference-in-means estimate ($93.36\pm 0.67\%$ of the absolute bias of an unrestricted design at $w_{Interference}=0.75$).
From the above, we reason that, when using a a difference-in-means estimator, it is beneficial to include mirror allocations and to not over-weight the interference metric, e.g. above 0.5. If using the regression estimator, however, mirror allocations can be excluded, and a larger $w_{Interference}$ can be applied without needing to worry about the variance of the estimator. We do note, however, that a difference-in-means estimator is generally less amenable to p-hacking, which is one factor to consider when choosing between these two estimators.
Figure 4: Bias, variance, and rejection rate of the difference-in-means estimator in an IGR design versus an unrestricted complete randomization (CR) design when mirror allocations are excluded (top) and when mirror allocations are included (bottom). Bias is plotted as a percentage relative to the bias of a CR design.
[…]